The no. of electrons present in 100ml of 0.1N solution of orthophosphoric acid??

Grade:12

1 Answers

Bhavya
7 years ago
Dear student
Electrons In 1 molecule of H3PO4 = 50
Thus 1 mole contains = 50 * NA electrons (NA = Avogadro’s Number)
Moles of acid in question = N*V = 100*0.1 = 10*10^-3 equivalents
Molarity = Normality / n-factor = 10^-2 /3
Electrons in these many moles of acid = 10^-2 * 50 * NA / 3